OpenStudy (anonymous):

$u^{2}_{x}/2−u_{y} = −x^{2}/2 ;\\ u(x, 0) = x$ How do we show that that the solution blows up in finite time and explain this in terms of characteristic of the equation.

OpenStudy (perl):

hmm, that equation is difficult to read

OpenStudy (michele_laino):

\[\begin{gathered} u_x^2/2 - {u_y} = - {x^2}/2; \hfill \\ u(x,0) = x \hfill \\ \end{gathered} \] I have rewrote them @perl